OpenStudy (howard-wolowitz):

I need a explanation to these:

OpenStudy (howard-wolowitz):

For the first one: A(x-5)+1(x-4)=-x+6 A(x-5)+x-4=-x+6 A(x-5)=-2x+10 now plug in x=6

OpenStudy (anonymous):

expain what? i.e. what are you trying to do? is this partial fraction decomposition?

OpenStudy (howard-wolowitz):

I know how to work these two! i got as far as that ^

OpenStudy (howard-wolowitz):

*dont know how

OpenStudy (howard-wolowitz):

and yes these are partial functions... if thats what their called

OpenStudy (anonymous):

you got this far \[A(x-5)=-2x+10\] right?

OpenStudy (howard-wolowitz):

yes

OpenStudy (anonymous):

you could a) distribute on the right and get \(Ax-10=-2x+10\) which means \(A=-2\) be equating the coefficients

OpenStudy (anonymous):

or you could b) factor on the left and get \[A(x-5)=-2(x-5)\] from which you see \(A=-2\) with your eyeballs

OpenStudy (anonymous):

lol i meant "distribute on the left" or "factore on the right" got those backwards sorry b

OpenStudy (howard-wolowitz):

I do see with my eyeballs. lol.. and I"ll try both ways. Thanks!

OpenStudy (anonymous):

damn i am full of typos if you distribute on the LEFT you get \[Ax-5A=-2x+10\]

OpenStudy (anonymous):

in any case it should be clear that \(A=-2\)

OpenStudy (howard-wolowitz):

I gotcha, thanks you for explaining that